# Polytope Volume Monitoring Problem: Formulation and Solution via Parametric Linear Program Based Control Barrier Function

### v1

@@ IV-C Synthesizing a NCBF-based Safety Filter
$M(z^{*}(x),x)=\operatorname{co}\left(\{{\mu}^{k}(x)\mid k\in\mathcal{I}_{V_{M}}(x)\}\right),$ (33)
@@
The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, i.e., the NCBF condition in (4) holds for $h_{r}(\cdot)$, if there exist $\alpha,\epsilon>0$, for all $x\in\mathcal{D}_{\epsilon}$, $\exists u\in\Psi(x)$, such that
@@
where $\Gamma(\mu,x,u):=-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)$ for any $\mu$.
@@
From the expression of $\partial_{x}L(\cdot)$ in (29), one can know that the scalar function $\partial_{x}L(\mu,z^{*},x)\left(f+gu\right)$ is a linear function of $\mu,z^{*},u$ respectively, hence the minimization optimization about $\mu$ in (32) is a LP at every $x$. Recalling the fact in [25, Sec 2.5, Theorem and Corollary 2] that the minimum value of a LP is obtained at the vertexes of the polytope, then
$\min_{\mu\in M(z^{*}(x),x)}\Gamma(\mu,x,u)$ $=\min_{\mu\in V_{M}(x)}\Gamma(\mu,x,u)$
$=\min_{{k}\in\mathcal{I}_{V_{M}}(x)}\Gamma(\mu^{k},x,u),$
@@
[…]
@@
where $\bar{u}^{*}(x):=\operatorname{col}(u^{*}(x),\delta^{*}(x))$, $u_{0}$ is any given continuous nominal control law, and the slack variable $\delta$ is introduced to guarantee the feasibility of such QP. $Q\in\mathbb{R}^{m\times m},\gamma>0$ are the (positive-definite) weighting matrix and scalar. The structure of the proposed safety filter (35) is summarized in Fig. 2.
Figure 2: A summary of the structure of the proposed safety filter (35).
@@
Consider the designed safety filter in the form of (35). For any system trajectory $x(t)$ starting from $x_{0}\in\mathcal{S}=\{x:h_{r}(x)\geq 0\}$, if
$u^{*}(x(t))\text{ is continuous, and }\delta^{*}(x(t))\equiv 0,$ (36)
then $x(t)\in\mathcal{S},\forall t\geq 0$, which implies that the PVM problem is solved. As a consequence,
$\Phi(x(t))\neq\emptyset,~{}u^{*}(x(t))\in\Psi(x(t)),\quad\forall t\geq 0,$ (37)
i.e., the output set-valued mapping never be empty and meanwhile the input constraint is never violated.
@@
Since Eq. (35b) denotes the soft constraints, the feasibility of the QP-based safety filter is guaranteed. The solution of closed-loop system is well-defined by the assumption that $u^{*}(x(t))$ is continuous. Meanwhile, the input constraint is never violated due to the hard constraints in (35c).
@@
$h^{\prime}_{r}(x(t);\dot{x}(t))\geq-\alpha h_{r}(x(t))-\delta^{*}(x(t)),$ (38)
where $\dot{x}(t)=f_{0}(x(t),u^{*}(x(t)))$. When $\delta^{*}(x(t))\equiv 0$, the analysis is obvious. In fact, from (38), $h^{\prime}_{r}(x(t);\dot{x}(t))\geq-\alpha h_{r}(x(t))$, by referring to Lemma 1, $h_{r}(x(t))\geq 0$, i.e., $x(t)\in\mathcal{S},\forall t\geq 0$. Recalling (12), the volume of $\Phi(x(t))$ is always greater than a threshold $\epsilon$. ∎
@@
In Theorem 3, two conditions in (36) are attached for the sake of completeness. They could be further relaxed. i) Known from (14), the inequality in (38) can be equivalently rewritten as:
${r^{*}}^{\prime}(x(t);\dot{x}(t))\geq-\alpha{r^{*}}(x(t))+{\alpha\epsilon_{0}-\delta^{*}(x(t))}.$ (39)
This reveals that if
$\|\delta^{*}(x(t))\|_{\mathcal{L}_{\infty}}:=\sup_{t}\|\delta^{*}(x(t)\|<\alpha\epsilon_{0},$
then ${r^{*}}^{\prime}(x(t);\dot{x}(t))>-\alpha{r^{*}}(x(t))$, which means that
${r^{*}}(x(t))>0,\forall t\geq 0.$
@@ V Numerical Simulation Example
Variables $x=[{p}_{x},{p}_{y},v,\theta]^{\top}$ are states, and $u=[a,\omega]^{\top}$ are control inputs, restricted in $\mathcal{U}=[0,2]\times[0,2]$. The reference controller in [9], with parameters $k_{p},k_{\theta},k_{v}$, is applied as the nominal control $u_{0}$ to drive the vehicle from the initial point to the goal. The barrier function candidate is designed as
@@
where $[{o}_{xi},{o}_{yi}]^{\top}$ is the center of the $i$-th obstacle and $R$ is the pre-defined safety distance. The positions of obstacles, initial and goal points can be seen from Fig. 3. The HOCBF [26] (with order $2$ and parameters $\alpha_{1}=10.0,\alpha_{2}=6.0$) is utilized to transform the position constraints $h_{i}(x)\geq 0$ to the affine constraints on input $u$, denoted as $u\in\mathcal{U}_{c}(x)$. According to Example 1, $\Phi(x)=\Psi(x)=\mathcal{U}\cap\mathcal{U}_{c}(x(t))$, the proposed safety filter (35) can be applied to solve the PVM problem here.
@@
(c) The time responses of $\delta^{*}(x(t))$ and $u^{*}(x(t))$.
@@
Fig. 3(a) shows the trajectories of the vehicle in the $XY$ plane. It can be observed that under the proposed safety filter (35), the vehicle successfully avoids the obstacles and reaches the goal in all cases. Under the method being compared (40), in Case 2, the vehicle can successfully avoid the obstacles and reaches the goal, while in Case 1 and Case 3, the vehicle would get stuck in the obstacles. Fig. 3(b) illustrates the time response of $r^{*}(x(t))$, which demonstrates the range of the feasible space of these two different safety filters. From it, one can observe that $r^{*}(x(t))$ is always greater than zero and greater than $\epsilon_{0}=0.6$ in most cases, which shows that the proposed method can maintain feasibility. As a comparison, the traditional one (40) becomes infeasible in Case 1 and Case 3, and its feasible space is much smaller than the proposed one’s in Case 2. In Fig. 3(c), the time responses of slack variable $\delta^{*}(x(t))$ and inputs $u^{*}(x(t))=[a(x(t)),\omega(x(t))]^{\top}$ under the proposed safety filter (35) in the above three cases are shown. One can observe that in all cases, $\delta^{*}(x(t))$ is equal to 00 approximately and much lower than the threshold $\alpha\epsilon_{0}=9$. Besides, the inputs $u^{*}(x(t))$ are continuous in all cases. These observations confirm the reasonability of the theoretical assumptions, analysis and conclusions in Theorem 3 and Remark 4.
@@ VI Conclusion
This paper formulates a class of control problems, called polytope volume monitoring (PVM), for which a parametric linear program-based control barrier function approach is proposed. This is achieved by establishing a bridge between the existing nonsmooth CBF and parametric optimization theory, where some useful lemmas about the continuity and directional differentiability of parametric programs are utilized to build nonsmooth CBF conditions, based on which a slacked QP-based safety filter is proposed to tackle the PVM problem step-by-step. The biggest limitation of this work is that the continuity of the proposed safety filter is not discussed formally. Future work involves designing continuous safety filters to tackle PVM problems, extending the proposed analysis and synthesis framework to other relevant problems and so on.
-A Nonsmooth Parametric Optimization
Let $\mathbb{T}\subset\mathbb{R}^{n}$ denote some open parameter set. For all $\bar{x}\in\mathbb{T}$, consider the convex parametric program with only inequalities:
$v(\bar{x}):=\min_{\bar{z}}$ $~{}\tilde{f}(\bar{z},\bar{x}),$ (A.1)
s.t. $\bar{z}\in{F}(\bar{x}):=\left\{\bar{z}\in\mathbb{R}^{n_{z}}\mid\tilde{g}_{j}(\bar{z},\bar{x})\leq 0,j\in\mathcal{J}\right\},$
with $\mathcal{J}:=\{1,2,\ldots,N\}$. For convenience, let the set of (global) minimizers according to (A.1) be denoted as
$S(\bar{x}):=\left\{\bar{z}\in F(\bar{x})\mid\tilde{f}(\bar{z},\bar{x})=v(\bar{x})\right\}.$ (A.2)
For $\bar{x}\in\mathbb{T}$ and feasible $\bar{z}\in F(\bar{x})$, the active index set is defined as $\mathcal{J}_{0}(\bar{z},\bar{x}):=\left\{j\in\mathcal{J}\mid\tilde{g}_{j}(\bar{z},\bar{x})=0\right\}$. Throughout this section, $\tilde{f},\tilde{g}_{j},j\in\mathcal{J}$, are continuously differentiable on $\mathbb{R}^{n_{z}}\times\mathbb{T}$, notation $\tilde{f},\tilde{g}_{j}\in\mathcal{C}^{1}$. The following Constraint Qualification conditions can be found in [18, (1.3) and (1.4)] and [27, (11) and (12)].
Definition 3.
@@
where ${\mu}:=\left({\mu}_{j},j\in\mathcal{J}_{0}(\bar{z},\bar{x})\right)\in\mathbb{R}^{\left|\mathcal{J}_{0}(\bar{z},\bar{x})\right|}$, collecting all multiplers $\mu_{j}$ corresponding to active constraints in $\mathcal{J}_{0}(\bar{z},\bar{x})$. Then the partial derivatives of $L(\cdot)$ are respectively given by
$\partial_{x}L(\mu,z,x):=$ $\partial_{x}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{x}\tilde{g}_{j}(z,x),$
$\partial_{z}L(\mu,z,x):=$ $\partial_{z}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{z}\tilde{g}_{j}(z,x).$
Lemma 3.
[18, Theorem 7.3] Consider the parametric convex program in (A.1) and let $F(\bar{x})$ be compact. If the MFCQ condition holds at $(\bar{z},\bar{x})$ satisfying $\bar{z}\in F(\bar{x})$, then $v(\cdot)$ is directional differentiable at $\bar{x}$ in any direction $d,\|d\|=1$ with
@@
where the minimizer set $S(\cdot)$ is defined in (A.2), and
$M(\bar{z},\bar{x})$ :=$:=\left\{\mu\mid\partial_{z}L(\mu,\bar{z},\bar{x})=0^{\top},\mu\geq 0\right\}.$ (A.6)
If LICQ holds, then $M(\bar{z},\bar{x})$ will be a singleton, i.e., the multiplier ${\mu}$ will be uniquely determined.
The min-max expression of directional derivative in (A.5) has also been pointed out by other literature [28, Theorem 2.4], [29, Theorem 4.24] yet under different conditions.
-B Proofs
Proposition 1 can be proved by some standard results in LP theory. Proposition 2 is the corollary of Lemma 3 with the boundness of set $M(z,x)$ to be proven.
Proof of Proposition 1: i) First, it is shown that $F_{P}(x)\neq\emptyset$ for all $x$. Recalling
$\Phi(x)=\{c\mid a_{j}^{\top}(x)c\leq b_{j}(x),~{}\forall j\in\mathcal{J}\},$
is nonempty, compact and convex, one can find a sufficiently small scalar $\varepsilon_{0}(x)>0$ and a point $c_{0}(x)$, such that
$a_{j}^{\top}(x)c_{0}(x)+\varepsilon_{0}(x)\leq b_{j}(x),~{}\forall j=1,2,\cdots,N.$
For every $j$, since $\|a_{j}(x)\|\neq 0$, there exists $c_{2}(x)\neq 0$, s.t.
$\varepsilon_{0}=a_{j}^{\top}c_{2}=\|a_{j}\|\underbrace{\|c_{2}\|\cos\langle a_{j},c_{2}\rangle}_{=:r_{j}(x)}>0.$
Defining $r_{0}:=\min_{j\in\mathcal{J}}~{}r_{j}$, one has $r_{0}>0$ and $\|a_{j}\|r_{0}\leq\varepsilon_{0}$, which leads to
$a_{j}^{\top}(x)c_{0}(x)+\|a_{j}(x)\|r_{0}(x)\leq b_{j}(x),~{}\forall j=1,2,\cdots,N,$
i.e., $F_{P}(x)\neq\emptyset$. Besides, $\Phi(x)$ is uniformly bounded for all $x$ as in Assumption 1, thus $F_{P}(x)$ is also uniformly bounded.
ii) Since the feasible set $F_{P}(x)$ is uniformly bounded, $r^{*}(x)$ and $h_{r}(x)$ are also uniformly bounded. Due to the fact that (14) is a LP problem, the optimization value is obtained at a point $z$ only if $z$ is on the boundary $\partial F_{P}(x)$, hence $S(x)\subset\partial F_{P}(x)$ is also uniformly bounded.
[…]
Proof of Proposition 2: According to (A.3), the local Lagrangian function near $\bar{x},\bar{z}$ is defined as:
$=\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial{{b}}_{j}}{\partial x}(x)\right),$
and $M(z,x):=\left\{\mu\mid\partial_{z}L(\mu,z,x)=0^{\top},~{}\mu\geq 0\right\}.$ In the above, the partial derivative of $L(\mu,z,x)$ w.r.t. variable $z,x$ at the point $\bar{z},\bar{x}$ are calculated. By further choosing $\bar{z}=z,\bar{x}=x$, the expressions in (29) and (30) can be obtained.
Since MFCQ condition holds and $F_{P}(x)$ is compact as proved in Proposition 1, then the conditions of Lemma 3 hold, from which one has
$h^{\prime}_{r}(x;\dot{x})$ $=-v^{\prime}_{r}(x;\dot{x})$
$=-\min_{z\in S(x)}\max_{\mu\in M(z,x)}\partial_{x}L(\mu,z,x)\dot{x}$
$=\max_{z\in S(x)}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\dot{x}.$
Hence, the expression in (28) is obtained.
The set $M(z,x)$ in (30) can be further rewritten as:
$M(z,x)$ =$=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\tilde{a}_{j}(x)\mu_{j}=-q,~{}\mu\geq 0\right\}$
$=M_{1}(z,x)\cap M_{2}(z,x),$
where
$M_{1}(z,x)$ :=$:=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}a_{j}(x)\mu_{j}=0\right\},$
$M_{2}(z,x)$ :=$:=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\|a_{j}(x)\|\mu_{j}=1,~{}\mu\geq 0\right\}.$
[…]

### v3

@@ IV-C Synthesizing a NCBF-based Safety Filter
$M(z^{*}(x),x)=\operatorname{co}\left(V_{M}(x)\right),$ (33)
@@
The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, if there exist $\alpha,\bar{\epsilon}>0$, for all $x\in\mathcal{D}_{\bar{\epsilon}}\mathrel{\mathop{\ordinarycolon}}=\{x\mid h_{r}(x)\geq-\bar{\epsilon}\}$, $\exists u\in\Psi(x)$, such that
@@
where $\Gamma(\mu,x,u)\mathrel{\mathop{\ordinarycolon}}=-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)$ for any $\mu$.
@@
From the expression of $\partial_{x}L(\cdot)$ in (29), one can know that the scalar function $\partial_{x}L(\mu,z^{*},x)\left(f+gu\right)$ is a linear function of $\mu,z^{*},u$ respectively, hence the minimization optimization about $\mu$ in (32) is a LP at every $x$. Recalling the fact in [26, Sec 2.5, Theorem and Corollary 2] that the minimum value of a LP is obtained at the vertexes of the polytope, then
$\min_{\mu\in M(z^{*}(x),x)}\Gamma(\mu,x,u)$ $=\min_{{k}\in\mathcal{I}_{V_{M}}(x)}\Gamma(\mu^{k},x,u),$
@@
$\text{s.t. }\Gamma(\mu^{k}(x),x,u)\geq-\alpha h_{r}(x)-\delta,~\forall k\in\mathcal{I}_{V_{M}}(x),$ (35b)
@@
where $\bar{u}^{*}(x)\mathrel{\mathop{\ordinarycolon}}=\operatorname{col}(u^{*}(x),\delta^{*}(x))$, $u_{0}$ is any given continuous nominal control law, and the slack variable $\delta$ is introduced to guarantee the feasibility of such QP. $Q\in\mathbb{R}^{m\times m},\gamma>0$ are the (positive-definite) weighting matrix and scalar respectively, where the penalty scalar $\gamma$ is much larger than the maximal eigenvalue of $Q$. The structure of the proposed safety filter (35) is summarized in Fig. 1 (b).
@@
If for the designed control law $\bar{u}^{*}(x)=\operatorname{col}(u^{*}(x),\delta^{*}(x))$ in (35), it holds at every $x\in\mathcal{S}=\{x\mid h_{r}(x)\geq 0\}$ that ${u}^{*}(x)$ is locally bounded and measurable, and the slack variable $\delta^{*}(x)$ is bounded by
$\delta^{*}(x)\leq\alpha\epsilon_{0}.$ (36)
Then the set $\mathcal{S}$ is forward invariant. As a consequence, for any trajectory $x(t)$ starting from $x_{0}\in\mathcal{S}$, it holds that
$x(t)\in\mathcal{A},~u^{*}(x(t))\in\Psi(x(t)),\quad\forall t\geq 0,$
i.e., the output set is always nonempty and input constraint is never violated.
@@
Since Eq. (35b) denotes the soft constraints, the feasibility of the QP-based safety filter is guaranteed whenever $\Psi(x)\neq\emptyset$. The solution of closed-loop system is well-defined by the assumption that ${u}^{*}(x(t))$ is be measurable and locally bounded [19]. Meanwhile, the input constraint is never violated due to the hard constraints in (35c).
@@
$(r^{*})^{\prime}(x(t);\dot{x}(t))$ $=h^{\prime}_{r}(x(t);\dot{x}(t))$ (37)
$\geq-\alpha h_{r}(x(t))-\delta^{*}(x(t))$
$=-\alpha r^{*}(x(t))+\alpha\epsilon_{0}-\delta^{*}(x(t)).$
When the inequality (36) holds, (37) will be reduced to $(r^{*})^{\prime}(x(t);\dot{x}(t))\geq-\alpha r^{*}(x(t))$. By referring to Lemma 1, the set $\mathcal{S}$ is forward invariant. Recalling the relationship in (11), one has $x(t)\in\mathcal{A}=\{x\mid\operatorname{Int}(\Phi(x))\neq\emptyset\},~\forall t\geq 0$. ∎
@@ V Numerical Simulation Example
Variables $x=[{p}_{x},{p}_{y},v,\theta]^{\top}$ are states, and $u=[a,\omega]^{\top}$ are control inputs, restricted in $\mathcal{U}=[-2,2]\times[-2,2]$. The reference controller in [9], with parameters $k_{p},k_{\theta},k_{v}$, is applied as the nominal control $u_{0}$ to drive the vehicle from the initial point to the goal. The barrier function candidate is designed as
@@
where ${o}_{xi},{o}_{yi}]^{\top}$ is the center of the $i$-th obstacle and $R$ is the pre-defined safety distance. The positions of obstacles, initial and goal points can be seen from Fig. 2. The HOCBF [28] (with order $2$ and parameters $\alpha_{1}=10.0,\alpha_{2}=6.0$) is utilized to transform the position constraints $h_{i}(x)\geq 0$ to the affine constraints on input $u$, denoted as $u\in\mathcal{U}_{c}(x)$. According to Example 1, $\Phi(x)=\Psi(x)=\mathcal{U}\cap\mathcal{U}_{c}(x(t))$, the proposed safety filter (35) can be applied to solve the PVM problem here.
@@
(c) Time responses of $\delta^{*}(x(t))$ and $u^{*}(x(t))$.
@@
Fig. 2(a) shows the trajectories of the vehicle in the $XY$ plane. It can be observed that under the proposed safety filter (35), the vehicle successfully avoids the obstacles and reaches the goal in all cases. Under the compared method (38), in Case 2, the vehicle can successfully avoid the obstacles and reaches the goal, while in Case 1 and Case 3, the vehicle would get stuck in the obstacles. Fig. 2(b) illustrates the time response of $r^{*}(x(t))$, which demonstrates the range of the feasible space of these two different safety filters. From it, one can observe that $r^{*}(x(t))$ is always greater than zero and greater than $\epsilon_{0}=0.6$ in most cases, which shows that the proposed method can maintain feasibility. As a comparison, the traditional one (38) becomes infeasible in Case 1 and Case 3, and its feasible space is much smaller than the proposed one’s in Case 2. In Fig. 2(c), the time responses of slack variable $\delta^{*}(x(t))$ and inputs $u^{*}(x(t))=[a(x(t)),\omega(x(t))]^{\top}$ under the proposed safety filter (35) in the above three cases are shown. One can observe that in all cases, $\delta^{*}(x(t))$ is equal to 0 approximately and much lower than the threshold $\alpha\epsilon_{0}=9$. Besides, the inputs $u^{*}(x(t))$ are continuous in all cases. These observations confirm the reasonability of the theoretical assumptions, analysis and conclusions in Theorem 3 and Remark 4.
@@ VI Conclusion
This paper formulates a class of control problems, called polytope volume monitoring (PVM), for which a parametric linear program-based control barrier function approach is proposed. This is achieved by establishing a bridge between the existing nonsmooth CBF and parametric optimization theory, where some useful lemmas about the continuity and directional differentiability of parametric programs are utilized to build nonsmooth CBF conditions, based on which a slacked QP-based safety filter is proposed to tackle the PVM problem step-by-step. The biggest limitation of this work is that the continuity of the proposed safety filter is not discussed formally. Future works involve designing continuous and smooth safety filters to tackle PVM problems, extending the proposed analysis and synthesis framework to other relevant problems and so on.
Appendix A Appendix
A-A Nonsmooth Parametric Optimization
Let $\mathbb{T}\subset\mathbb{R}^{n}$ denote some open parameter set. For all $\bar{x}\in\mathbb{T}$, consider the convex parametric program:
$v(\bar{x})\mathrel{\mathop{\ordinarycolon}}=\min_{\bar{z}}$ $~\tilde{f}(\bar{z},\bar{x}),$ (A.1)
s.t. $\bar{z}\in{F}(\bar{x})\mathrel{\mathop{\ordinarycolon}}=\left\{\bar{z}\in\mathbb{R}^{n_{z}}\mid\tilde{g}_{j}(\bar{z},\bar{x})\leq 0,j\in\mathcal{J}\right\},$
with $\mathcal{J}\mathrel{\mathop{\ordinarycolon}}=\{1,2,\ldots,N\}$. For convenience, let the set of (global) minimizers according to (A.1) be denoted as
$S(\bar{x})\mathrel{\mathop{\ordinarycolon}}=\left\{\bar{z}\in F(\bar{x})\mid\tilde{f}(\bar{z},\bar{x})=v(\bar{x})\right\}.$ (A.2)
For $\bar{x}\in\mathbb{T}$ and feasible $\bar{z}\in F(\bar{x})$, the active index set is defined as $\mathcal{J}_{0}(\bar{z},\bar{x})\mathrel{\mathop{\ordinarycolon}}=\left\{j\in\mathcal{J}\mid\tilde{g}_{j}(\bar{z},\bar{x})=0\right\}$. Throughout this section, $\tilde{f},\tilde{g}_{j},j\in\mathcal{J}$, are continuously differentiable on $\mathbb{R}^{n_{z}}\times\mathbb{T}$, notation $\tilde{f},\tilde{g}_{j}\in\mathcal{C}^{1}$. The following Constraint Qualification conditions can be found in [18, (1.3) and (1.4)] and [23, (11) and (12)].
Definition A.1.
@@ A-A Nonsmooth Parametric Optimization
where ${\mu}\mathrel{\mathop{\ordinarycolon}}=\left({\mu}_{j},j\in\mathcal{J}_{0}(\bar{z},\bar{x})\right)\in\mathbb{R}^{\left|\mathcal{J}_{0}(\bar{z},\bar{x})\right|}$, collecting all multiplers $\mu_{j}$ corresponding to active constraints in $\mathcal{J}_{0}(\bar{z},\bar{x})$. Then the partial derivatives of $L(\cdot)$ are respectively given by
$\partial_{x}L(\mu,z,x)\mathrel{\mathop{\ordinarycolon}}$= $\partial_{x}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{x}\tilde{g}_{j}(z,x),$
$\partial_{z}L(\mu,z,x)\mathrel{\mathop{\ordinarycolon}}$= $\partial_{z}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{z}\tilde{g}_{j}(z,x).$
Lemma A.2.
[18, Theorem 7.3] Consider the parametric convex program in (A.1) and let $F(\bar{x})$ be compact. If MFCQ holds at $(\bar{z},\bar{x})$ satisfying $\bar{z}\in F(\bar{x})$, then $v(\cdot)$ is directional differentiable at $\bar{x}$ in any direction $d,\|d\|=1$ with
@@
where the minimizer set $S(\cdot)$ is defined in (A.2), and the set of multipliers $M(\cdot)$ is defined as
$M(\bar{z},\bar{x})$ $\mathrel{\mathop{\ordinarycolon}}=\left\{\mu\mid\partial_{z}L(\mu,\bar{z},\bar{x})=0^{\top},\mu\geq 0\right\}.$ (A.6)
If LICQ holds, then $M(\bar{z},\bar{x})$ will be a singleton, i.e., the multiplier $\mu$} will be uniquely determined [18, EX. 2.6]. The min-max expression of directional derivative in (A.5) has also been pointed out by other literature [29, Theorem 2.4], [30, Theorem 4.24] yet under different conditions.